### 10 Mar 2024 v2

Abstract
In the abstract Tile Assembly Model, self-assembling systems consisting of tiles of different colors can form structures on which colored patterns are “painted.” We explore the complexity, in terms of the numbers of unique tile types required, of assembling various patterns.We first demonstrate how to efficiently self-assemble a set of simple patterns, then show tight bounds on the tile type complexity of self-assembling 2-colored patterns on the surfaces of square assemblies. Finally, we demonstrate an exponential gap in tile type complexity of self-assembling an infinite series of patterns between systems restricted to one plane versus those allowed two planes.
1 Introduction
During the process of self-assembly, a disorganized collection of components experiencing only random motion and local interactions combine to form structures. Examples of self-assembly abound in nature, from crystals to cellular components, and these systems have inspired researchers to study them to better understand the underlying principles governing them as well as to engineer systems that mimic them. Along the spectrum of both natural and artificial self-assembling systems are those which (1) use very small numbers of component types and form simple, unbounded, repeating patterns, (2) use a large number of component types, on the order of the entire sizes of the structures created, to form structures with highly-specified, asymmetric designs, and (3) those which use very small numbers of components to make arbitrarily large, bounded or unbounded, symmetric or asymmetric structures whose growth is directed algorithmically.
Systems in category (2), so-called fully-addressed or hard-coded have the benefit of being able to uniquely define each “pixel” of the structure and therefore to “paint” arbitrary pictures, which we’ll refer to as patterns, on their surfaces[21, 23, 17]. Although it is easy to see that in models of tile-based self-assembly such as the abstract Tile Assembly Model (aTAM)[22] any finite structure or pattern can self-assemble from a hard-coded system, it has previously been shown that there exist infinite structures and patterns that cannot self-assemble in the aTAM [12, 13]. The benefits of the algorithmic self-assembly[24, 8, 18] of category (3) include precise formation of shapes using exponentially fewer types of components[19, 20], thus reducing the cost, reducing the effort to fabricate and implement, and increasing the speed of growth of these systems[5]. However, the drastic reduction in the number of component types means a corresponding increase in their reuse. This results in copies of the same component type appearing in many locations throughout the resulting target structure, thus removing the ability to uniquely address each pixel when forming patterns. This generally results in a reduction in the number of patterns that are producible.
In this paper we study the trade-off between the numbers of unique components, or tile types, needed to self-assemble designed patterns and the complexities of the patterns that can self-assemble. Past lines of work have dealt with the self-assembly of patterns in the aTAM from the perspective of the computational complexity of designing minimal tile sets to self-assemble given patterns (the so-called “PATS” problem)[11, 4, 15, 14] and others have shown the possibilities and impossibilities of assembling some classes of infinite patterns[12, 16]. In contrast, our first results present constructions for making tile sets that self-assemble a series of relatively simple patterns to demonstrate how efficiently they can be built algorithmically. These consist of patterns of white and black pixels on the surfaces of squares, and include (1) a pattern with a single black pixel, (2) a pattern with some number $k$ of black pixels, and (3) grids of alternating black and white stripes. All of these are shown to have exponential reductions in tile type requirements, a.k.a. tile complexity, over fully-addressed systems.
Our next pair of results combine to show a tight bound on the tile complexity of self-assembling arbitrary patterns of two colors on the surfaces of $n\times n$ squares for almost all patterns. Using an information theoretic argument we prove a lower bound, namely that for almost all patterns on $n\times n$ squares, such patterns have tile complexity $\Omega\left(\frac{n^{2}}{\log{n}}\right)$. We then provide a construction that, when given an arbitrary $n\times n$ pattern of black and white pixels, generates a tile set of $O\left(\frac{n^{2}}{\log{n}}\right)$ tile types that self-assemble an $n\times n$ square with black and white tiles that form that pattern. Although this is not a significant improvement over the $n^{2}$ tile types required to naively implement a fully-addressed set of tile types, the lower bound proves that for almost all patterns this is the best possible tile complexity.
Although the prior result showed that any $n\times n$ pattern can self-assemble using $O\left(\frac{n^{2}}{\log{n}}\right)$ tile types, our next result shows that, if given two planes in which tiles can self-assemble (one on top of the other) then it is possible for some patterns to self-assemble using exponentially fewer tile types than when systems are restricted to a single plane. (In fact, this result can be modified for arbitrary separation in tile complexity.) The proof of this result uses a novel application of diagonalization to tile-based self-assembly. Namely, one system simulates every system of a given tile complexity class for a bounded number of steps, sequentially and within a square of one plane, in order to algorithmically generate a pattern that is guaranteed not to be made by any of those systems, and then prints that generated pattern on the square of the second plane above the assembly that performed the simulations. We also show how to extend these square patterns infinitely to cover the plane, while maintaining the same tile complexity argument.
Overall, our results demonstrate boundaries on tile complexities of algorithmic self-assembling systems when forming patterns, and help to demonstrate their benefits over fully-addressed systems. To make some of our results easier to understand, we have created a set of programs and tile sets that can be used to view examples. These can be found online: Pattern Self-Assembly Software [7]. Due to space limitations, proofs have been moved to a Technical Appendix in this version, and this full version can also be found online [6].
2 Preliminary Definitions and Models
In this section we define the terminology and model used throughout the paper.
2.1 The abstract Tile-Assembly Model
We work within the abstract Tile-Assembly Model[22] in 2 and 3 dimensions. We will use the abbreviation aTAM to refer to the 2D model, 3DaTAM for the 3D model, and barely-3DaTAM to refer to the 3D model when restricted to the use of only 2 planes of the third dimension (a.k.a. the “just barely 3D aTAM”), meaning that tiles can only be placed in locations with $z$ coordinates equal to 00 or $1$ (use of the other two dimensions is unbounded). These definitions are borrowed from [9] and we note that [19] and [13] are good introductions to the model for unfamiliar readers.
Let $\mathbb{N}$ be the set of nonnegative integers, and for $n\in\mathbb{N}$, let $[n]=\{0,1,...,n-2,n-1\}$. Fix $d\in\{2,3\}$ to be the number of dimensions and $\Sigma$ to be some alphabet with $\Sigma^{*}$ its finite strings. A glue $g\in\Sigma^{*}\times\mathbb{N}$ consists of a finite string label and non-negative integer strength. There is a single glue of strength 00, referred to as the null glue. A tile type is a tuple $t\in(\Sigma^{*}\times\mathbb{N})^{2d}$, thought of as a unit square or cube with a glue on each side. A tile set is a finite set of tile types. We always assume a finite set of tile types, but allow an infinite number of copies of each tile type to occupy locations in the $\mathbb{Z}^{d}$ lattice, each called a tile.
Given a tile set $T$, a configuration is an arrangement (possibly empty) of tiles in the lattice $\mathbb{Z}^{d}$, i.e. a partial function $\alpha:\mathbb{Z}^{d}\dashrightarrow T$. Two adjacent tiles in a configuration interact, or are bound or attached, if the glues on their abutting sides are equal (in both label and strength) and have positive strength. Each configuration $\alpha$ induces a binding graph $B_{\alpha}$ whose vertices are those points occupied by tiles, with an edge of weight $s$ between two vertices if the corresponding tiles interact with strength $s$. An assembly is a configuration whose domain (as a graph) is connected and non-empty. The shape $S_{\alpha}\subseteq\mathbb{Z}^{d}$ of assembly $\alpha$ is the domain of $\alpha$. For some $\tau\in\mathbb{Z}^{+}$, an assembly $\alpha$ is $\tau$-stable if every cut of $B_{\alpha}$ has weight at least $\tau$, i.e. a $\tau$-stable assembly cannot be split into two pieces without separating bound tiles whose shared glues have cumulative strength $\tau$. Given two assemblies $\alpha,\beta$, we say $\alpha$ is a subassembly of $\beta$ (denoted $\alpha\sqsubseteq\beta$) if $S_{\alpha}\subseteq S_{\beta}$ and for all $p\in S_{\alpha}$, $\alpha(p)=\beta(p)$ (i.e., they have tiles of the same types in all locations of $\alpha$).
A tile-assembly system (TAS) is a triple $\mathcal{T}=(T,\sigma,\tau)$, where $T$ is a tile set, $\sigma$ is a finite $\tau$-stable assembly called the seed assembly, and $\tau\in\mathbb{Z}^{+}$ is called the binding threshold (a.k.a. temperature). If the seed $\sigma$ consists of a single tile, i.e. $|\sigma|=1$, we say $\mathcal{T}$ is singly-seeded. Given a TAS $\mathcal{T}=(T,\sigma,\tau)$ and two $\tau$-stable assemblies $\alpha$ and $\beta$, we say that $\alpha$ $\mathcal{T}$-produces $\beta$ in one step (written $\alpha\to^{\mathcal{T}}_{1}\beta$) if $\alpha\sqsubseteq\beta$ and $|S_{\beta}\setminus S_{\alpha}|=1$. That is, $\alpha\to^{\mathcal{T}}_{1}\beta$ if $\beta$ differs from $\alpha$ by the addition of a single tile. The $\mathcal{T}$-frontier is the set $\partial^{\mathcal{T}}\alpha=\bigcup_{\alpha\to^{\mathcal{T}}_{1}\beta}S_{\beta}\setminus S_{\alpha}$ of locations in which a tile could $\tau$-stably attach to $\alpha$. When $\mathcal{T}$ is clear from context we simply refer to these as the frontier locations.
We use $\mathcal{A}^{T}$ to denote the set of all assemblies of tiles in tile set $T$. Given a TAS $\mathcal{T}=(T,\sigma,\tau)$, a sequence of $k\in\mathbb{Z}^{+}\cup\{\infty\}$ assemblies $\alpha_{0},\alpha_{1},\ldots$ over $\mathcal{A}^{T}$ is called a $\mathcal{T}$-assembly sequence if, for all $1\leq i<k$, $\alpha_{i-1}\to^{\mathcal{T}}_{1}\alpha_{i}$. The result of an assembly sequence is the unique limiting assembly of the sequence. For finite assembly sequences, this is the final assembly; whereas for infinite assembly sequences, this is the assembly consisting of all tiles from any assembly in the sequence. We say that $\alpha$ $\mathcal{T}$-produces $\beta$ (denoted $\alpha\to^{\mathcal{T}}\beta$) if there is a $\mathcal{T}$-assembly sequence starting with $\alpha$ whose result is $\beta$. We say $\alpha$ is $\mathcal{T}$-producible if $\sigma\to^{\mathcal{T}}\alpha$ and write $\mathcal{A}[\mathcal{\mathcal{T}}]$ to denote the set of $\mathcal{T}$-producible assemblies. We say $\alpha$ is $\mathcal{T}$-terminal if $\alpha$ is $\tau$-stable and there exists no assembly that is $\mathcal{T}$-producible from $\alpha$. We denote the set of $\mathcal{T}$-producible and $\mathcal{T}$-terminal assemblies by $\mathcal{A}_{\Box}[\mathcal{\mathcal{T}}]$. If $|\mathcal{A}_{\Box}[\mathcal{\mathcal{T}}]|=1$, i.e., there is exactly one terminal assembly, we say that $\mathcal{T}$ is directed. When $\mathcal{T}$ is clear from context, we may omit $\mathcal{T}$ from notation.
2.2 Patterns
Let $C$ be a set of colors and let $P\subseteq(\mathbb{Z}^{d}\times C)$. We say that $P$ is a $d$-dimensional pattern, i.e., a set of locations and corresponding colors. Let ${\rm dom}\;P$ be the set of locations that are assigned a color. A pattern is k-colored when the number of unique colors used is $k$. Let $\texttt{Color}(P,\vec{l})$ be a function that takes a pattern and a location $\vec{l}$ and returns the color of the pattern at that location. (and is undefined if $\vec{l}\not\in{\rm dom}\;(P)$).
Given a TAS $\mathcal{T}=(T,\sigma,\tau)$, we allow each tile type to be assigned exactly one color from some set of colors $C$. Let $C_{P}\subseteq C$ be a subset of those colors, and $T_{C_{P}}\subseteq T$ be the subset of tiles of $T$ whose colors are in $C_{P}$. Given an assembly $\alpha\in\mathcal{A}[\mathcal{\mathcal{T}}]$, we use ${\rm dom}\;(\alpha)$ to denote the set of all locations with tiles in $\alpha$ and ${\rm dom}\;_{C_{p}}(\alpha)$ to denote the set of all locations of tiles in $\alpha$ with colors in $C_{P}$. Given a location $\vec{l}\in\mathbb{Z}^{d}$, let $\texttt{Color}(\alpha,\vec{l})$ define a function that takes as input an assembly and a location and returns the color of the tile at that location (and is undefined if $\vec{l}\not\in{\rm dom}\;(\alpha)$). We say $\mathcal{T}$ weakly self-assembles pattern $P$ iff for all $\alpha\in\mathcal{A}_{\Box}[\mathcal{T}]$, ${\rm dom}\;_{C_{P}}(\alpha)=P$ and $\forall(\vec{l},c)\in P,c=\texttt{Color}(\alpha,\vec{l})$. We say $\mathcal{T}$ strictly self-assembles pattern $P$ iff $T_{C_{P}}=T$, i.e. all tiles of $T$ are colored from $C_{P}$, and $\mathcal{T}$ weakly self-assembles $P$ (i.e. all locations receiving tiles are within $P$).
Let the set of all patterns be $\mathbb{P}$, the set of all $c$-colored patterns be $\mathbb{P}_{c}$, and $\texttt{SQPATS}_{c,n}\subset\mathbb{P}_{c}$ be the set of all $c$-colored patterns that are on the surfaces of $n\times n$ squares. Let $\texttt{SQPATS}_{c}=\bigcup_{n\in\mathbb{Z}^{+}}{\texttt{SQPATS}_{c,n}}$, i.e., the set of all $c$-colored patterns that are on the surfaces of squares. A pattern class is an infinite set of patterns parameterized by some set of values $X$, and can be represented as a function $PC:X\rightarrow\mathbb{P}$ that maps parameters $X$ to some pattern $P\in\mathbb{P}$. Let $\mathbb{T}$ be the set of all aTAM systems. A construction for pattern class $PC$ is a function $C_{PC}:\mathbb{P}\rightarrow\mathbb{T}$ that takes a pattern $P\in\mathbb{P}$ and outputs an aTAM system $\mathcal{T}\in\mathbb{T}$ such that $\mathcal{T}$ weakly self-assembles $P$.
3 Simple Patterns
In this section, we define several relatively simple pattern classes and present constructions that can build them efficiently.
First we define a pattern class whose patterns each consist of a 2-colored $n\times n$ square that is completely white except for a single black pixel.
Definition 1 (Single-Pixel Pattern Class)
Given $n,i,j\in\mathbb{N}$, where $i,j<n$, define $\texttt{SinglePixel}(n,i,j)\rightarrow P$ such that (1) $P\in\texttt{SQPATS}_{2,n}$, (2) $\forall\vec{l}\in dom(P)-\{(i,j)\}$, $\texttt{C}olor(P,\vec{l})=\text{White}$, and (3) $\texttt{C}olor(P,(i,j))=\text{Black}$.
Theorem 3.1
For all $n,i,j\in\mathbb{N}$ such that $n\geq i,j$, there exists an aTAM system $\mathcal{T}=(T,\sigma,2)$ such that $|\sigma|=1$, $|T|=O(\log(n))$, and $\mathcal{T}$ weakly self-assembles $\texttt{SinglePixel}(n,i,j)$.
Proof
We present a construction that, given an $n,i,j\in\mathbb{N}$ such that $i,j<n$, creates a TAS $\mathcal{T}$ with tile complexity of $O(\log n)$ that weakly self-assembles $P=\texttt{SinglePixel}(n,i,j)$. Figure 1 shows a high-level depiction of how we build the assembly.
The assembly starts by growing a hard-coded rectangle with an empty interior called a counter box. Each side of the box has glues for a counter to attach. The side lengths are $s=\lceil\log n\rceil$, as this is the maximum length needed to encode the bits required for a binary counter to count a full dimension of the entire square (which is the maximum that could be necessary). Thus, the counter box uses $O(\log n)$ tiles types.
Each side of the counter box has a number encoded in the outward-facing glues. These numbers are pre-computed so that binary counter tiles (i.e. sets of tiles that operate as standard binary counters) grow outward from them to form a cross-like structure that extends to each boundary of the $n\times n$ square. A constant-sized set of tile types (independent of $i$,$j$, and $n$) is used for each of the four counters. To complete the $n\times n$ square, a constant-sized set of filler tiles fill in between the counters and inside the counter box.
The seed tile is a single black tile, from which the counter box grows, while all other tile types are white. If the black tile is within $\log n$ of the side of the square, then the counter box grows away from the edge, meaning the counter box always has room to grow inside of the $n\times n$ square.
We have shown that the tile complexity of this assembly is $O(\log n)$ and that it weakly self-assembles $P$. Thus, the Theorem 3.1 is proved.
Figure 1: (a) An example of a single-pixel pattern. The black pixel is located at (10, 2). (b) The same single-pixel pattern but with the counter box and counter tiles colored for demonstration. The counter box is colored red. The counters are colored blue. The white locations are filled by generic filler tiles.
Next, we define a pattern class whose patterns each consist of a 2-colored $n\times n$ square that is completely white except for a set of (separated and individual) black pixels.
Definition 2 (Multi-Pixel Pattern Class)
Given $n\in\mathbb{N}$ and a set of locations $L\subseteq[0,n-1]^{2}$ such that $\forall(x,y),(x^{\prime},y^{\prime})\in L$, $|x-x^{\prime}|\geq\lceil\log n\rceil\lor|y-y^{\prime}|\geq\lceil\log n\rceil$, define $\texttt{MultiPixel}(n,L)\rightarrow P$ such that (1) $P\in\texttt{SQPATS}_{2,n}$, and (2) $\forall\vec{v}\in dom(P),\texttt{C}olor(P,\vec{v})=$ Black if $\vec{v}\in L$ else White.
Figure 2: (a) An example of a multi-pixel pattern with three black pixels. (b) A tree of counters is constructed to grow to each pixel and the edges of the square. The counter boxes are colored red. The counters are colored blue. The white locations are filled by generic filler tiles.
Theorem 3.2
For all $n\in\mathbb{N}$ and sets of locations $L\subseteq[0,n-1]^{2}$ such that $\forall(x,y),(x^{\prime},y^{\prime})\in L$, $|x-x^{\prime}|\geq\lceil\log n\rceil\lor|y-y^{\prime}|\geq\lceil\log n\rceil$, there exists an aTAM system $\mathcal{T}=(T,\sigma,2)$ such that $|\sigma|=1$, $|T|=O(|L|\log n)$, and $\mathcal{T}$ weakly self-assembles $\texttt{MultiPixel}(n,L)$.
The proof of Theorem 3.2 uses a construction that is an extension of that used in the proof of Theorem 3.1, and essentially uses a series of counters to build a path connecting all pixels and filler tiles for the remaining portion of the square. A high-level depiction is shown in Figure 2, and full details can be found in Section 7.1.
The final (relatively) simple pattern class that we define contains patterns that each consist of a $2$-colored $n\times n$ square with a set of repeating black horizontal rows and a set of repeating black vertical columns.
Definition 3 (Stripes Pattern Class)
Given $n,i,j\in\mathbb{N}$, where $i,j<n$, define $\texttt{Stripes}(n,i,j)\rightarrow P$ such that (1) $P\in\texttt{SQPATS}_{2,n}$, and (2) $\forall x,y\in[0,n-1],\texttt{C}olor(P,(x,y))=\text{Black}\text{ if }x\mod i=0\text{ or }y\mod j=0,\text{ else White}$.
Theorem 3.3
For all $n,i,j\in\mathbb{N}$, where $i,j<n$, there exists an aTAM system $\mathcal{T}=(T,\sigma,2)$ such that $|\sigma|=1$, $|T|=O(\log(n))$ and $\mathcal{T}$ weakly self-assembles $\texttt{Stripes}(n,i,j)$.
The proof of Theorem 3.3 can be found in Section 7.2. It is done by construction, where the construction has counters that grow vertically to count to, and mark, the locations of horizontal strips, and counters that grow horizontally to count to, and mark, the locations of vertical stripes. Counters also keep track of the distance to the boundaries of the square to ensure growth stops at the correct locations. An overview can be seen in Figure 3.
Figure 3: (a) An example of a stripes pattern. (b) The blue tiles count to the next stripe, while the red tiles count the number of stripes. Green tiles represent the starting rows for the counters (with the seed tile being at the corner where the green row and column intersect). Dark grey tiles represent counter tiles that are colored black
4 Tight Bounds for Patterns on $n\times n$ Squares
In this section, we prove tight bounds on the tile complexity of self-assembling 2-colored patterns on the surfaces of $n\times n$ squares for almost all such patterns.
Theorem 4.1
For almost all positive integers $n$ and $P\in\texttt{SQPATS}_{2,n}$, the tile complexity of weakly self-assembling $P$ by a singly-seeded system in the aTAM is $\Theta\left(\frac{n^{2}}{\log n}\right)$.
We prove Theorem 4.1 by separately proving the lower and upper bounds, as Lemma 1 and Lemma 2, respectively.
Lemma 1
For almost all patterns $P\in\texttt{SQPATS}_{2}$, the tile complexity of weakly self-assembling $P$ by a singly-seeded system in the aTAM is $\Omega\left(\frac{n^{2}}{\log n}\right)$.
The proof of Lemma 1 is a straight-forward information-theoretic argument and can be found in Section 8.
To prove the upper bound for Theorem 4.1, we prove the following, which is a stronger result that applies to all positive integers $n$.
(a) The skeleton. The seed is represented in green in the lower left and the arrows show the directions of growth.
(b) The square once the ribs of the skeleton have filled in (blue growing to the left, yellow growing to the right).
Figure 4: A schematic example of the construction of the proof of Lemma 2. Instead of showing the black and white colors corresponding to the pattern, we color the tiles to show the pieces of the construction to which they belong.
Lemma 2
For all positive integers $n$ and $P\in\texttt{SQPATS}_{2,n}$, there exists an aTAM system $\mathcal{T}=(T,\sigma,1)$ such that $|\sigma|=1$, $|T|=O\left(\frac{n^{2}}{\log n}\right)$ and $\mathcal{T}$ weakly self-assembles $P$.
Proof
We proceed by construction. Let $n\in\mathbb{Z}^{+}$ be the dimensions of the square and $P$ be the $n\times n$ pattern of black and white pixels to weakly self-assemble on the square. Our construction will yield a system $\mathcal{T}=(T,\sigma,1)$ that self-assembles an $n\times n$ square on which $P$ is formed by the black and white tiles of $T$. The tile set $T$ will be composed of two subsets, $T_{s}$ whose tiles form the skeleton, and $T_{r}$ whose tiles form the ribs. We first explain the formation of the skeleton, then that of the ribs. Figure 4 shows a high-level depiction.
Skeleton The seed is part of the skeleton and is placed at location $(\lfloor\log n\rfloor,0)$ and is given the color $\texttt{Color}(P,(\lfloor\log n\rfloor,0))$. Since a vertical column of the skeleton has width one, and the ribs growing off of each side have length $\lfloor\log n\rfloor$, the width of a pair of ribs and its skeleton column (which we will call a rib-pair) is $12\lfloor\log n\rfloor+$ 1. Dividing the full width $n$ by the width of a rib-pair, and taking the floor, gives the number of full rib-pairs that will fit. Let $f=\lfloor\frac{n}{2\log n+1}\rfloor$ be this number. Let $r=n\mod(2\lfloor\log n\rfloor+1)$ be the remaining width after the last full rib-pair. If $r<\lfloor\log n\rfloor+1$, then a column of the skeleton grows up immediately to the right of the last full rib-pair, and its ribs are of length $r-1$ and grow to the right. If $r\geq\lfloor\log n\rfloor+1$, then the last skeleton column grows upward $\lfloor\log n\rfloor$ positions to the right of the last full rib-pair and has full-length ribs (i.e., $\lfloor\log n\rfloor$) that grow to its left and ribs of length $r-(\lfloor\log n\rfloor+1)$ grow to its right. In the first case, the row of the skeleton that forms the bottom row of the square extends from the seed to $x$-coordinate $f(2\lfloor\log n\rfloor+1)+1$. In the second case, that row extends from the seed to $x$-coordinate $f(2\lfloor\log n\rfloor+1)+\lfloor\log n\rfloor+1$. The tiles of that row are hard-coded and there are $O(n)$ of them. Starting with the seed and then occurring at every $12\lfloor\log n\rfloor+$ 1 locations of the bottom row, a hard-coded set of tiles grows a column of height $n-1$. This row and set of columns are the full skeleton. The number of tile types is $O(n)$ for the row and $O(n)$ for each of the $O\left(\frac{n}{\log n}\right)$ columns, for a total of $O(n)+O\left(\frac{n^{2}}{\log n}\right)=O\left(\frac{n^{2}}{\log n}\right)$ tile types. Note that each skeleton tile type is given the color of the corresponding location in the pattern $P$.
Ribs From the east and west sides of each location on the columns of the skeleton, ribs grow. Each rib is composed of $\lfloor\log n\rfloor$ tiles (except the ribs growing from the easternmost column, which may be shorter). Since there are two possible colors for each of the $\lfloor\log n\rfloor$ locations of a rib, there are a maximum of $2^{\lfloor\log n\rfloor}\leq n$ possible color patterns for any rib to match the corresponding locations in $P$. (Note that we will discuss the construction of the tiles for ribs that grow to the east, and for ribs that grow to the west the directions are simply reversed.) For any given rib $r$, let the portion of $P$ corresponding to the locations of $r$ be represented by the binary string of length $\lfloor\log n\rfloor$ where each black location is represented by a 00, and each white by a $1$. For example, for a rib $r$ of length $5$ growing eastward from a column, if the corresponding locations of $P$ are “black, black, white, black, white”, then the binary string will be “$00101$”. For each possible binary string $b$ of length $\lfloor\log n\rfloor$, i.e. $b\in\{0,1\}^{\lfloor\log n\rfloor}$, a unique tile type, $t_{b}$, is made. with the glue $b$ on its west side and the glue $b[1:]$ (i.e. $b$ with its left bit truncated) on its east side. This tile type is given the color corresponding to the first bit of $b$. Additionally, for each skeleton column tile from which a rib should grow to the east with pattern $b$, the glue $b$ will be on its east side, allowing $t_{b}$ to attach. This results in the creation of a maximum of $n$ unique tile types (and there will be another $n$ for the first tiles of each westward growing rib). Recall that the tile types for the skeleton were already accounted for and each is hard-coded so that the placement of these glues does not require any new tile types for the skeleton.
Now, the process is repeated for each binary string from length $b-1$ to $1$, with the color of each tile being set to the value of the first remaining bit. Each iteration requires half as many tile types to be created as the previous, i.e. $2^{\lfloor\log n\rfloor-1}$, then $2^{\lfloor\log n\rfloor-2}$, $\ldots$, $2$. Intuitively, each rib position has glues that encode their bit value in the pattern and the portion of the pattern that must be extended outward from them, away from the skeleton. Therefore, for the last position on the tip of each rib, there are exactly 2 choices, white or black, and so all ribs share from a set of two tile types made specially for the ends of ribs. For the tile types of ribs that grow to the east, the total summation is $\Sigma^{\log n-1}_{x=0}{2^{\log n-x}}=2n-2=O(n)$. Accounting for the additional tile types needed for westward growing ribs, the full tile complexity of the ribs is $O(n)$.
Thus, the total tile complexity for the tile types of the skeleton plus those of the ribs is $O\left(\frac{n^{2}}{\log n}\right)+O(n)=O\left(\frac{n^{2}}{\log n}\right)$.
Correctness of construction The system $\mathcal{T}$ designed to weakly self-assemble $P$, as discussed, has a seed of a single tile, and since all tile attachments require forming a bond with a single neighbor, the temperature of the system can be $\tau=1$. Our prior analysis shows that the tile complexity is correct at $O\left(\frac{n^{2}}{\log n}\right)$, and showing that $\mathcal{T}$ weakly self-assembles $P$ is trivial since (1) the tiles of the skeleton are specifically hard-coded to be colored for their corresponding locations in $P$, and (2) for each possible pattern corresponding to a rib there is a hard-coded set of rib tiles that match that pattern and grow from the skeleton into those locations. Thus, $P$ is formed and Lemma 2 is proved, and with both Lemmas 1 and 2, Theorem 4.1 is proved. (Example aTAM systems for this construction, as well as software capable of generating other systems for patterns derived from image files, and for simulating them, can be found online [7].)
5 Repeated Patterns
Figure 5: An example of an assembly that repeats a pattern $m=5$ times horizontally and vertically. Each spine is colored solely for clarity of presentation, and in the actual construction, the colors of the tiles on the spines would match the pixels of the pattern. Red spines represent a 1, and blue spines represent a 0. The spines count upwards until the counter is finished
In this section, we discuss patterns consisting of repeated, arbitrary square sub-patterns, and that efficient systems exist that weakly self-assemble them.
Definition 4 (Grid Repeat Pattern Class)
Given $n,m\in\mathbb{Z}^{+}$ and $P\in\texttt{SQPATS}_{2,n}$, define $\texttt{GridRepeat}(P,m)\rightarrow P^{\prime}$ such that $P^{\prime}\in\texttt{SQPATS}_{2,nm}$ is an $nm\times nm$ square consisting of an $m\times m$ square composed of an $n\times n$ grid of copies of the pattern $P$.
Theorem 5.1 (Repeated Pattern Tile Complexity)
For all $n,m\in\mathbb{N}$, and $P\in\texttt{SQPATS}_{2,n}$ there exists an aTAM system $\mathcal{T}=(T,\sigma,2)$ such that $|\sigma|=1$, $|T|=O(\frac{n^{2}}{\log{n}}+\log{mn})$ and $\mathcal{T}$ weakly self-assembles $\texttt{GridRepeat}(P,m)$.
The proof of Theorem 5.1 can be found in Section 9. It makes use of an extension of the construction for the proof of Lemma 2 and embeds a counter into the skeleton and ribs so that the copies of the sub-pattern are correctly counted.
6 Barely-3DaTAM patterns
In this section, we show that there exist patterns, both finite and infinite, that can be weakly self-assembled using exponentially fewer tile types by barely-3DaTAM systems than by any regular, 2D aTAM systems. (We also note that the exponential separation can be increased arbitrarily.)
Theorem 6.1
For all $n\in\mathbb{Z}^{+}$, for some $m\in\mathbb{Z}^{+}$ there exists a 7-colored $m\times m$ pattern, $p_{n}$, such that (1) no aTAM system $\mathcal{T}_{\leq n}=(T,\sigma,\tau)$ exists where $|T|\leq n$, $|\sigma|=1$, and $\mathcal{T}_{\leq n}$ weakly self-assembles $p_{n}$, but (2) a barely-3DaTAM system $\mathcal{T}_{p_{n}}=(T_{p_{n}},\sigma_{p_{n}},2)$ exists where $|T_{p_{n}}|=O(\log{n}/\log{\log{n}})$, $|\sigma_{p_{n}}|=1$, and $\mathcal{T}_{p_{n}}$ weakly self-assembles $p_{n}$.
(a) Example grid pattern for bit sequence $11010101$.
(b) Example grid pattern for bit sequence $00101010$.
Figure 6: Example $p_{n}$ patterns created by the construction in the proof of Theorem 6.1. The (repeatedly copied) binary sequence derived from the results of the simulations of aTAM systems starts at the top, with the two colors of that row, and all subsequent boundary rows, being determined by the first bit of that pattern. During the downward growth from that row, during which the full $m\times m$ square is formed, the repeating grid formed by the copies of that pattern is copied both downward and to both sides. The boundary columns also have two colors determined by the first bit of the sequence (one of them the same as in the boundary rows) for a total of 3 boundary colors. The interiors always use the same 4 colors.
Proof sketch
[…]
The bit sequence that determines the colors used by the boundaries, and the ordering of the colors on the boundaries and in the interiors, is determined via simulations of a series of aTAM systems. Intuitively, our proof utilizes a construction that performs a diagonalization against all possible aTAM systems with $\leq n$ tile types by simulating each for a bounded number of steps, and for each keeping track of the color of tile it places in a location specific to the index of that system so that it can ultimately generate the colored pattern $p_{n}$ that differs in at least one location from every simulated system.
Figure 7: Schematic overview of the portion of the construction for the proof of Theorem 6.1 that grows in plane $z=0$. Modules are not shown to scale. (Green) The seed tile, (Fuchsia) the base conversion module that unpacks the binary representation of $n$, (Aqua) the module that simulates the Turing machine $M$ on input $n$, which itself simulates each aTAM system with $\leq n$ tile types in sequence and saves a result bit (Black or White) for each, (Yellow and Grey) the pattern of result bits is copied repetitively to the right until it covers the entire top row. (Light Grey) A “filler” tile causes the assembly to form a complete square.
The dimensions of each $c\times c$ cell are $c=\texttt{SF}(n)$, where $\texttt{SF}(n)$ is a function that takes a number of tile types and returns an upper bound on the number of all possible singly-seeded aTAM systems with $\leq n$ tile types and $\leq 8$ colors. (Note that $\texttt{SF}(n)$ is actually greater than the number of such systems, and details of $\texttt{SF}(n)$ can be found in Section 10.) The colors of the rows and columns encode the bit sequence generated by the simulations, with the same bit sequence encoded in both the rows and the columns via an assignment of colors. There is a unique color assigned for each intersection of two bits (i.e. 00, 01, 10, and 11), with 4 colors reserved for boundaries of grid cells and 4 separate colors reserved for the interior locations of the grid cells. Therefore, the colored pattern of every cell differs from the assemblies produced by all aTAM systems with $\leq n$ tile types. It forms on the top layer of a two-layered $m\times m$ square where $m=O(n^{21n})$, and the barely-3DaTAM system that forms it uses $O(\log{n}/\log{\log{n}})$ tiles since the tiles types for all modules are constant except those that encode $n$ using optimal encoding [1]. (Note that the value of $m$ could be smaller, $O(n^{4n}n^{8})$, if it wasn’t desired that both planes be the same size. Additionally, by having $M$ simulate systems with larger tile sets, the value of $m$ would increase but the difference in tile complexity between the barely-3DaTAM system and the systems incapable of making its patterns could be increased beyond the current exponential bound.) Additional technical details, including pseudocode for the algorithms of the Turing machine $M$ and its simulations of all systems with $\leq n$ tile types, the layout of data structures used during the simulation of a system, and time complexity analysis, can be found in Section 10 of the technical appendix.
6.1 Extending a pattern $p_{n}$ to infinitely cover $\mathbb{Z}^{2}$
Although it is already known that there are infinite patterns that can’t weakly self-assemble from any finite-sized tile set [12], the following corollary simply shows how the previously defined patterns can be extended to infinitely cover the plane, while keeping an arbitrary spread in the tile complexity required by aTAM and barely-3DaTAM systems.
Corollary 1
For all $n\in\mathbb{Z}^{+}$, there exists a 7-colored pattern, $p_{n_{\infty}}$, that infinitely covers the plane $\mathbb{Z}^{2}$ such that (1) no aTAM system $\mathcal{T}_{\leq n}=(T,\sigma,\tau)$ exists where $|T|\leq n$, $|\sigma|=1$, and $\mathcal{T}_{\leq n}$ weakly self-assembles $p_{n_{\infty}}$, but (2) a barely-3DaTAM system $\mathcal{T}_{p_{n}}=(T_{p_{n}},\sigma_{p_{n}},2)$ exists where $|T_{p_{n}}|=O(\log{n}/\log{\log{n}})$, $|\sigma_{p_{n}}|=1$, and $\mathcal{T}_{p_{n}}$ weakly self-assembles $p_{n_{\infty}}$.
To prove Corollary 1, we extend the construction from the proof of Theorem 6.1 so that every pattern $p_{n}$ from the proof of Theorem 6.1 is extended to infinitely cover the $\mathbb{Z}^{2}$ plane, becoming $p_{n_{\infty}}$, by usage of “grid-reconstruction,” i.e., a method of copying the square grid infinitely to each side. This requires $O(1)$ unique tile types in addition to those used in the previous construction. Due to the symmetry exhibited by all $m\times m$ squares of $p_{n}$ patterns along their northeast $\rightarrow$ southwest diagonal, copies of the same pattern may be copied along these diagonals infinitely. Details of the construction can be found in Section 11.